solve by completing the square
x^2 + 8x - 9 = 0

Answer by scott8148(6628)

x^2 + 8x = 9

adding half of the x-term coefficient, squared ___ x^2 + 8x + (8/2)^2 = 9 + (8/2)^2

x^2 + 8x + 16 = 25 ___ (x + 4)^2 = 5^2

taking square root ___ x + 4 = ±5

x + 4 = 5 ___ x = 1

x + 4 = -5 ___ x = -9